OpenStudy (anonymous):

Need help on solving three system of equations using substitution? -x-y-z=-9 -5x+5y+6z=8 3x+3z=5 I really need help...

OpenStudy (anonymous):

This is confusing....

OpenStudy (anonymous):

Somebody please help me...

OpenStudy (anonymous):

x-y-z=-9 z = -9 -x + y Using this: 3x+3z=5 3x + 3(-9 -x +y) = 5 3x -27 -3x +3y = 5 3y - 27 = 5 3y = 32 y = 32/3 x - 32/3 - (-9 - x + 32/3) = -9 x - 32/3 + 9 + x - 32/3 = -9 2x - 12.3 = -9 2x = 3.3 x = 1 2/3 z = -9 - 1 2/3 + 32/3 z = 0 1 2/3 - 32/3 - 0 = -9

OpenStudy (anonymous):

Thanks for helping

OpenStudy (anonymous):

R u sure these how you do them

OpenStudy (anonymous):

Best Response 0 x-y-z=-9 z = -9 -x + y Using this: 3x+3z=5 3x + 3(-9 -x +y) = 5 3x -27 -3x +3y = 5 3y - 27 = 5 3y = 32 y = 32/3 x - 32/3 - (-9 - x + 32/3) = -9 x - 32/3 + 9 + x - 32/3 = -9 2x - 12.3 = -9 2x = 3.3 x = 1 2/3 z = -9 - 1 2/3 + 32/3 z = 0 1 2/3 - 32/3 - 0 = -9

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

That was quick

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Can you add her people as friends

OpenStudy (anonymous):

here

OpenStudy (anonymous):

who you talking about

OpenStudy (anonymous):

Oh nothing..never mind

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

How did you get 12.3